OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 15x - 5y = 30. intercept = (-2, 0); y-intercept = (0, 6) intercept = (0, -6); y-intercept = (2, 0) x-intercept = (2, 0); y-intercept = (0, -6) intercept = (0, 6); y-intercept = (-2, 0) please help i think its C

OpenStudy (***[isuru]***):

hi, if the graph was in the shape of y = mx + c c is the "y" coordinate of "y - intercept".... so .. let's arrange the 15x - 5y = 30 in to that shape... 15x - 5y = 30 divide the whole by 5 3x - y = 6 y = 3x - 6 .... now the function is in the form of y = mx + c y = 3x - 6 so.. "y" coordinate of "y - intercept" is (-6) now the coordinate point of y intercept should be something like this (x , -6) the only answer which contain a (-6) as the "y" coordinate of "y - intercept" is 3rd one... so ... u can dump the other 3 answers easily.. ur answer then will be 3 hope this will help ya!!

OpenStudy (anonymous):

thank you so much so it was C

OpenStudy (***[isuru]***):

yep!!

OpenStudy (***[isuru]***):

u r welcome!!